OpenStudy (anonymous):

In a country club, 80% of women play tennis, 20% play pool and 10% play both tennis and pool. if a women is chosen at random, find the probability: a) that she plays tennis or pool b) that she plays neither tennis or pool. c) if she plays tennis what is the probability she plays pool? d) if she plays pool what is the probability she play tennis?

OpenStudy (anonymous):

A) p(A or B)=p(a)+p(b)-p(a and b) = (.8)+(.2)-(.1) = .9 = 90% B) p( not A or B) = 1-p(A or B) = 10% (the complement of above) C) 10%/80% = 1/8 D) 10%/20% = .5

OpenStudy (anonymous):

Thanks